OpenStudy (vshiroky):

If Professor Juarez had 30 students in her class, and she wanted to test her hypothesis using the 5% level of significance, what cutoff t score would she use? (You should be able to figure this out without a table because only one answer is in the correct region.) A) 304.11 B) 1.699 C) .113 D) 2.500

OpenStudy (vshiroky):

I think this is B if someone can double check.. @agent0smith

OpenStudy (vshiroky):

@kirbykirby

OpenStudy (kirbykirby):

Looks right to me. It's one-tailed?

OpenStudy (vshiroky):

It doesn't say.. I'm assuming so... would you be able to double check like 6 more for me? lol and I need help with 3 that I can't figure out.. Sorry to be a pest.. This class is killing me lol

OpenStudy (kirbykirby):

Even if it's two-tailed, it would be the answer that's still closest...

OpenStudy (vshiroky):

For the following question(s): A school counselor tests the level of depression in fourth graders in a particular class of 20 students. The counselor wants to know whether the kind of students in this class differs from that of fourth graders in general at her school. On the test, a score of 10 indicates severe depression, while a score of 0 indicates no depression. From reports, she is able to find out about past testing. Fourth graders at her school usually score 5 on the scale, but the variation is not known. Her sample of 20 fifth graders has a mean depression score of 4.4. Use the .01 level of significance. 1. The counselor calculates the unbiased estimate of the population’s variance to be 15. What is the variance of the distribution of means? A) 15/20 = 0.75 B) 15/19 = 0.79 C) 152/20 = 11.25 D) 152/19 = 11.84 2. Suppose the counselor tested the null hypothesis that fourth graders in this class were less depressed than those at the school generally. She figures her t score to be .20. What decision should she make regarding the null hypothesis? A) Reject it B) Fail to reject it C) Postpone any decisions until a more conclusive study could be conducted D) There is not enough information given to make a decision 3. Suppose the standard deviation she figures (the square root of the unbiased estimate of the population variance) is .85. What is the effect size? A) 5/.85 = 5.88 B) .85/5 = .17 C) (5  4.4)/.85 = .71 D) .85/(5  4.4) = 1.42

OpenStudy (kirbykirby):

Sure, if I can help

OpenStudy (vshiroky):

For those I got 1.a 2.b 3.c

OpenStudy (kirbykirby):

What are the squares? On my scree, like in question 2, it says "She figures her t score to be .20" and I see more squares in question 3

OpenStudy (kirbykirby):

screen*

OpenStudy (vshiroky):

negative

OpenStudy (vshiroky):

I'm not sure if you responded.. my open study started to be a little screwy

OpenStudy (kirbykirby):

I agree with 1 For 2, I would maybe agree with b, but it might be d since we don't know if the population is normally distributed (it's not mentioned... although it might be assumed). 3 looks good too. sorry I'm slow... I was trying to read up on effect size.. I'm not too familiar with it.

OpenStudy (vshiroky):

my kik took a poo earlier lol